Solve the following exponential equation

\left(\frac{5}{6}\right)^{3x+7}=\left(\frac{36}{25}\right)^{2x}

Please show every step, thank you!
Answer #1

3x+7 Glven (3) (36 25 this we solve applying log on both side by 3x+7 28 log o logom C3 25 log (F) 3x++ lorg ( mlogb] 3x+73x+7 = -48 7 =-42-32 7=-70 x= -21 t x=-) i Answer is x= -1